OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

parallel: same slope y=3x+b -2=3(4) + b -2=12+b -14=b y=3x-14